I thought I would indicate some of the changes which were implemented
recently at IOTA.

1) picomotor pulse control box, otherwise known as tick-tick stopper -
You can now limit which axis of the picomotor is active
with a small box connected between the Picomotor control box
and the hand paddle.  I only made one of these, i.e., for the
control box above the alignment telescope.  It has three
separate toggle switches to enable the A, B or C axis.  It has
a label to indicate which switch controls what, and which
position is enabled.  It has indicator lights: Green - safe
condition, you can walk away and the motors won't drift; Red -
ready to move axis. The lights are not integral to the circuit
but just indicators for the user, if the batteries die and the
lights go out, the box will still work.

2) New Shutters -
Solenoid activated shutters had a tendency to stick when left
on (i.e., closed) for a while.  The new system uses a
mechanism similar to the flip in LED lights with DC motors and
on/off limit switches.  It operates at 5V and 15V.   The inrush
& stall current on the new motors is greater than the power supply
rating - if this becomes a problem and we need to change power supply
you may want to swap to the old shutters temporarily.  The old
shutters will remain in place and you can switch back in a few
minutes if necessary.
Just be careful handling the 12V wires on the old system.
I will post instructions on the web site soon.

3) Connections for feed mirrors -
Because the wires to the system enter near the 5m port, there
are wires which may need to be changed when working at 0m, 5m
and >5m for the A and C telescopes.  I will write up
instructions soon or change the wiring a bit to make it more

4) Polarized mode -
The Wollaston prism can be inserted between the two input
lenses of the Picnic and the two polarizations can be seen on
the detector.  Guy Perrin believes that the accuracy is improved
if the data was taken in each of the polarizations.  The
fibers apparently change their birefringence with small
changes in temperature which leads to mixing of the light from
different polarization states from the different telescopes.

5) New Control system -
ICS 4 is about ready to be used.  It currently has the same
functionality as the previous system, and data has been
collected with it.  There were a few items which needed fixing
so it was not ready on October 1.  The new control system is
based on a new and improved framework and hopefully will work
better.  Database recording of configurations is implemented.

As new features (not quite ready to be used), the new control
system allows access to the AAIB (Auto Alignment Interface
Box) which allows you to do remote alignment (not auto yet,
see next item).  AAIB also allows you to change the filter
without going into the other room.  (The positioning is not as
accurate as it should be so this feature is still being worked

Starting ICS 4 has slightly different commands, but is still as
straight forward as in the previous system. Instructions are
on site and will also be available online.

6) Remote Alignment system -
We have a beamsplitter on a slider which can be inserted
remotely into any of the three beams.  The slider can be
controlled from the Dell PC in the observing area (the PC near
the star tracker) and will soon be incorporated in ICS-4.  The
beamsplitter feeds a camera with a lens whose focus, iris and
zoom are controlled from the AAIB box.  The alignment
(flip-in) lights at the exit, Short delay, Long delay, corner,
feed, Piezo block and Secondary can all be controlled from
this box in ICS4.  The box also interfaces to the pointing of
the A and B feed mirrors, the yaw corrections and the three
piezo blocks.

The remote alignment system also enables you to move all of
the picomotors in the system: The alignment beam splitter, the
3 exit mirrors, the 2 SD yaws, the 3 corner mirrors, the C
Feed mirror, the 3 star tracker mirrors, the 3 Picnic feed

Using the AAIB system will entail an hour moving the
connections and has not been fully tested on the mirrors in
the system yet.

Note on Fall 2005 startup -
or what delayed implementing more changes in 5) and 6)

The Forest Fire in July forced me off the mountain for a couple of
weeks and the unplugging was not quite as thorough as last year.  The
motor controller for the long delay line got blown out by lightning
and took 10 days debug and repair.  The motor power supply for LD2
failed and was replaced.  Some of the home and limit switches on the
long delay were also blown and replaced.  One of the interface chips
to the C telescope limits went into intermittent failure mode and was
replaced.  The Temperature controller for the Star tracker camera had
a blown component which required sending it back to the
manufacturer; it is now back and operational.  The Technobox Altera
card for the Picnic camera went into a more permanent failure mode and
was replaced with a spare; the original went back to the manufacturer
for evaluation.  Thunder storms continued through mid September so
reconnecting equipment and testing was delayed.